# Relative pose of three calibrated and partially calibrated cameras from four points using virtual correspondences

### v2

@@ 3.1 Calibrated cameras
We use a lightweight architecture with a backbone of shared MLP layers, similar to [10], such that each triplet of correspondences is processed independently. The input to our network is a $4\times 6$ matrix of 4 point correspondences containing the $x$ and $y$ coordinates in three views. In estimating the epipolar geometry, the order of the point correspondences does not matter. Thus, our network is designed to be permutation invariant in that input axis. The input is normalized as follows: We apply a rotation matrix to the points in each view independently, so that the mean point of the first three points is sent to $(0,0)$, and the fourth point is sent to $(0,y)$. Let $\mathbf{m}^{1},\mathbf{m}^{2}$, and $\mathbf{m}^{3}$ be the mean points of the three corresponding points in three views. We aim to predict the corresponding point of $\mathbf{m}^{1}$ in the second view. Let us denote this predicted point by $\mathbf{\tilde{m}}^{2}$, i.e., our $5^{th}$ virtual correspondence in the first two views will be $\mathbf{m}^{1}\leftrightarrow\mathbf{\tilde{m}}^{2}$. As suggested by the 4p3v(M) solver, $\mathbf{\tilde{m}}^{2}$ should be close to $\mathbf{m}^{2}$. Thus, we use the mean point $\mathbf{m}^{2}$ as the initialization of our network and predict a shift from $\mathbf{m}^{2}$.
@@
We train and validate our network using synthetic data. The dataset is generated as follows: We generate a scene of 10K random 3D points. To generate each instance, we randomly select four 3D points and three cameras with random rotations and translations that see the scene, and project those points into the cameras. In total, we use $1$M such instances, which we split 70% for training and 30% for validation. The network was implemented in PyTorch [41] and trained for 30 epochs using the Adam optimizer [25]. More details can be found in the supp. material.
Similarly to the 4p3v(M) solver, the 4p3v(L) solver solves the problem using the efficient 5pt+P3P solver.
4p3v(L±$\pm\delta$) and 4p3v(L±$\pm\delta$init) solvers:  Similar to 4p3v(M±$\pm\delta$), we try to compensate for potential noise in the predictions $\tilde{\mathbf{m}}^{2}$ returned by the network by running the 5pt solver for the first two views three times for three different virtual correspondences. We test two variants: (1) In the 4p3v(L±$\pm\delta$) solver, we add a shift ±$\pm\delta$ to the predicted points $\tilde{\mathbf{m}}^{2}$, similar to the 4p3v(M±$\pm\delta$) solver. In this way, we generate two more virtual correspondences. (2) In the 4p3v(L±$\pm\delta$init) solver, we add a shift $\delta$ to the initialization $\mathbf{m}^{2}$ of the network. Thus, we run the network three times with the initializations $\mathbf{m}^{2}$ and two points $\mathbf{m}^{2}_{\delta}$ to predict three different virtual correspondences.
3.2 Partially calibrated cameras
Our novel solvers for the 4p3vf problem of estimating the relative pose of three cameras with an unknown shared focal length from four correspondences follow the idea of the virtual correspondences used in our 4p3v solvers for calibrated cameras. The idea is to transform the extremely complex 4p3vf problem into the 6pt+P3P problem for which efficient solutions exist. The 6pt+P3P solver first estimates the unknown focal length and the relative pose of the first two cameras using the efficient 6pt solver [59] and then registers the third camera using the P3P solver [43].
This means that, in contrast to the 4p3v solvers presented in Section 3.1 that generate only one virtual correspondence (+ potentially additional shifted versions of this correspondence), in our novel 4p3vf solvers, we need to generate two new virtual correspondences to obtain six correspondences in the first two views.
In the 4p3vf(M) solver, we generate these two new virtual correspondences using the mean points of two different triplets of corresponding points in the first two cameras.
In the 4p3vf(L) solver, we run the network described in Section 3.1 for two different initializations of mean points of different triplets in the first camera.
4p3vf(M±$\pm\delta$), 4p3vf(L±$\pm\delta$), 4p3vf(L±$\pm\delta$init) solvers work similarly to the calibrated $\delta$-solvers More details on all 4p3vf solvers can be found in the supp. mat.

### v3

@@ 8 L-based Solvers
We use a lightweight architecture with a backbone of shared MLP layers, similar to [6], so that each triplet of correspondences is processed independently. The input to our network is a $4\times 6$ matrix of 4 point correspondences where the $i^{th}$ row contains the $x$ and $y$ coordinates of the $i^{th}$ point in three views, i.e., the $i^{th}$ correspondence. In estimating the epipolar geometry, the order of the point correspondences does not matter. Thus, our network is designed to be permutation invariant on that input axis. The input is normalized as follows: We apply a rotation matrix to the points in each view independently, so that the mean point of the first three points is sent to $(0,0)$, and the fourth point is sent to $(0,y)$. Let $\mathbf{m}^{1},\mathbf{m}^{2}$, and $\mathbf{m}^{3}$ be the mean points of the three corresponding points in three views. We aim to predict the corresponding point of $\mathbf{m}^{1}$ in the second view. Let us denote this predicted point by $\mathbf{\tilde{m}}^{2}$, i.e., our $5^{th}$ virtual correspondence in the first two views will be $\mathbf{m}^{1}\leftrightarrow\mathbf{\tilde{m}}^{2}$. As suggested by the 4p3v(M) solver, $\mathbf{\tilde{m}}^{2}$ should be close to $\mathbf{m}^{2}$. Thus, we use the mean point $\mathbf{m}^{2}$ as the initialization of our network and predict a shift from $\mathbf{m}^{2}$.
We use a simple MLP-based backbone, similar to [6, 42]. More precisely, our input consists of four 6-vectors, of the $x,$ and $y$ coordinates of four point correspondences in three views. The first part of the model is a shared MLP 3-block of dimensions 6, 32, and 32, exporting a 32-dimensional feature for every correspondence. Then we apply a channel-wise max pooling aggregation, which is then concatenated at the end of each 32-dimensional feature. This results in having an 64-dimensional vector for each correspondence, which are passed into another shared MLP 3-block of 64, 64, and 64 dimensions. We aggregate those vectors via a max pooling function to get a 64-dimensional vector encoding, which eventually passes through an MLP 3-block to reduce the dimension gradually from 64, to 32, and finally to 2, which will be the prediction of the $x$ and $y$ coordinates in the second camera. In all MLPs, in the first 2 layers of the blocks, we use a Leaky RELU activation function [58] with slope 0.01. As for the last layer of the MLPs, in the first two blocks, we use a RELU activation function, while in the last MLP we use a $\tanh$ activation, since we want the output to be in the range $(-1,1)$. For a visualization of the architecture, see Figure 8.
We used a simple network architecture to show the promise of 4p3v(L)/4p3vf(L) solvers, namely that learning can produce more accurate point correspondences. The experiments shown in the paper verify this behavior. We believe that more advanced network architectures (e.g., using a set transformer architecture [29]) have the potential to improve the results even more and reduce the gap between the 4p3v/4p3vf solvers and the oracle 4p3v(O)/4p3vf(O) solvers.
@@
For both training and validation, we use synthetic data. Our synthetic dataset contains 1M input instances. 70% of the dataset is used for training while the remaining 30% is used for validating the performance of the network. We generate 10K 3D points inside a 10$\times$10$\times$10 cube, and to generate each instance, we pick 4 random points and project them to 3 cameras with random rotations and translations that view the cube from a random distance between 20 to 50 units.
4p3v(L±$\pm\delta$), 4p3vf(L±$\pm\delta$), and 4p3vf(L±$\pm\delta$init) solvers.  Similar to the 4p3v(M±$\pm\delta$) solver, we try to compensate for potential noise in the prediction $\tilde{\mathbf{m}}^{2}$ returned by the network by running the 5pt solver for the first two views three times for three different virtual correspondences. We test two variants: (1) In the 4p3v(L±$\pm\delta$) and 4p3vf(L±$\pm\delta$) solvers, we add a shift ±$\pm\delta$ to the predicted point $\tilde{\mathbf{m}}^{2}$, similar to the 4p3v(M±$\pm\delta$) solver. In this way, we generate two additional virtual correspondences. (2) In the 4p3v(L±$\pm\delta$init) solver, we add a shift $\delta$ to the initialization $\mathbf{m}^{2}$ of the network. Thus, we run the network three times with three initializations, namely $\mathbf{m}^{2}$, $\mathbf{m}^{2}_{+\delta}$, and $\mathbf{m}^{2}_{-\delta}$. Each initialization affects the normalization of the points in the 2nd view, in the sense that a different point (the initialization) is sent to (0, 0), leading to a different input, thus predicting different virtual correspondence.